A student leaves the house,26 km due to north, then turns onto a street and continues in a direction 30° NE for 35 km and finally turns onto the highway due east for 40 km. What is her total displacement from his house?


1
Expert's answer
2021-05-21T17:06:20-0400

"y=26+35\sin30°=43.5\ (km)"


"x=40+35\cos30°=70.31\ (km)"


"L=\sqrt{43.5^2+70.31^2}=82.68\ (km)" . Answer